# Critical Iterative Denoising A Discrete Generative Model Applied to Graphs

### v1

@@ Abstract
Discrete Diffusion and Flow Matching models have significantly advanced generative modeling for discrete structures, including graphs. However, the time dependencies in the noising process of these models lead to error accumulation and propagation during the backward process. This issue, particularly pronounced in mask diffusion, is a known limitation in sequence modeling and, as we demonstrate, also impacts discrete diffusion models for graphs.
To address this problem, we propose a novel framework called Iterative Denoising, which simplifies discrete diffusion and circumvents the issue by assuming conditional independence across time. Additionally, we enhance our model by incorporating a Critic, which during generation selectively retains or corrupts elements in an instance based on their likelihood under the data distribution.
Our empirical evaluations demonstrate that the proposed method significantly outperforms existing discrete diffusion baselines in graph generation tasks.
Machine Learning, ICML
@@ 1 Introduction
Denoising models, such as Discrete Diffusion and Discrete Flow Matching, have significantly advanced generative modeling for discrete structures (Austin et al., 2021; Campbell et al., 2022, 2024; Gat et al., 2024), including graphs (Haefeli et al., 2022; Vignac et al., 2023). Despite their success, these models suffer from an important limitation caused by the time dependency in their noising and denoising processes. Errors introduced early in the process accumulate and propagate, degrading generative performance. This issue is particularly pronounced and apparent in mask diffusion, where the noising process progressively masks elements of an instance. This challenge has been identified in discrete sequence modeling, and solutions such as corrector sampling have been proposed to mitigate it.
@@
Figure 1: Validity rate (without correction) of generated molecules trained on Zinc250k as a function of the Number of Function Evaluations (NFE) for three models: Discrete Diffusion (baseline), Iterative Denoising (ours), and Critical Iterative Denoising (ours).
@@
In graph modeling, the marginal distribution across node and edge attributes has proven to be a more suitable noise distribution than the mask distribution, as corroborated by our experiments. However, we show both theoretically and empirically that the issue of error propagation due to time dependencies also affects discrete diffusion and flow matching models in graph generation, where the noise corresponds to the marginal distributions.
@@
To address this challenge, we introduce a novel framework called Iterative Denoising, which simplifies discrete diffusion by assuming conditionally independent noise across time. By removing the direct dependency on partially denoised instances from the previous step our framework facilitates error correction and substantially improves generative performance.
@@
Furthermore, we show that our model can be interpreted as selectively corrupting certain elements while leaving others unmodified. Building on this insight, we introduce a Critic, which modulates the corruption probability of each element based on the element’s likelihood under the data distribution. We theoretically motivate our approach and we empirically show that the method further enhances models’ performances.
@@
Our empirical evaluations highlight the effectiveness of the proposed framework, demonstrating superior performance over existing discrete diffusion baselines in graph generation tasks. Notably, our method tackles the challenge of generating valid molecular structures, achieving reliable performance for molecules. In Figure 1 we see that our method needs only a small number of denoising steps to reach almost 100% molecule validity, while the standard Discrete Diffusion tops at 80%.
Our code will be publicly released upon acceptance and is available to reviewers upon request.
2 Background and Related Works
@@
Graph diffusion models are powerful methods for generating discrete graph structures, such as molecules. However, existing approaches, including Discrete Diffusion Models and Discrete Flow Matching (DFM), face challenges due to the temporal dependencies of the noising process. These dependencies introduce error propagation and accumulation, ultimately degrading generative performance. This section introduces core concepts, discusses these limitations, and motivates our approach.
@@ 2.1 Notation
We define a graph as a set of nodes and edges, denoted by ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$. A graph is represented by its adjacency matrix ${\bm{A}}\in[d_{A}+1]^{n\times n}$, where $n=|{\mathcal{V}}|$, and $[d_{A}+1]=\{1,...,d_{A}+1\}$, with $d_{A}$ representing the number of edge types, and the additional label corresponding to the absence of an edge. Node attributes, if present, are encoded as integers in an annotation vector ${\bm{x}}\in[d_{X}]^{n}$, where $d_{X}$ is the number of node types.
To facilitate readability and simplify the method’s transferability to other discrete modalities, such as sequence modeling, we use $z$ to denote a single element. In graph modeling, $z$ may refer to either a node or an edge attribute when the formulation applies to both $x^{(i)}$ and $a^{(i,j)}$. The corresponding capital symbol $Z$ represents the entire instance, such as a graph ${\mathcal{G}}$, z. For simplicity and consistency with common practice, we use $p(x)$ to denote the Probability Mass Function (PMF) $P(X=x)$. The distribution of a data point element is represented via the Dirac delta distribution, denoted as $\delta_{z_{1}}(z)$, with all mass concentrated on $z_{1}$. We sometimes represent the univariate categorical distribution over the variable $p(z)$ as a vector ${\bm{z}}$, where the $i$-th component $z_{i}$ denotes the probability that $z$ belongs to the category indexed by $i$.
We refer to the mask distribution as the distribution where all probability mass is concentrated on an additional synthetic attribute labeled as Mask, denoted by $\delta_{\text{Mask}}(z)$. This distribution is sometimes referred to as the absorbing-state distribution (Austin et al., 2021). We define mask diffusion as the class of diffusion models that employ this distribution as noise.
We use $q$ to refer to the noising distributions and $p$ for the denoising ones.
@@ 2.2 Discrete Diffusion Models
where $\tau=\{1-\Delta_{t},...,t+\Delta_{t},t\}$.
The continuous case is described by taking $\lim_{\Delta_{t}\to 0}q_{t|1}(z\mid z_{1})$, which transforms the product into a geometric integral. In discrete time, a forward diffusion step can be represented with transition probability matrices: $[{\bm{Q}}_{t}]_{i,j}=q(z_{t}=j\mid z_{t+\Delta_{t}}=i)$ (Vignac et al., 2023). A diffusion step then consists of sampling from:
$q_{{t-\Delta_{t}}\mid t}(z\mid z_{t})=\text{Cat}(z;{\bm{z}}_{t}^{T}{\bm{Q}}_{t-\Delta_{t}}).$
Similarly, we have:
$q_{t\mid{1}}(z\mid z_{1})=\text{Cat}(z;{\bm{z}}_{1}^{T}\bar{{\bm{Q}}}_{t}),$
where $\bar{{\bm{Q}}}_{t}=\prod_{r\in\tau}{\bm{Q}}_{r}$.
In this work, we focus on transition matrices of the form:
${\bm{Q}}_{t}=(1-\beta_{t}){\bm{I}}+\beta_{t}{\bm{N}},$ (2)
where ${\bm{I}}$ is the identity matrix and ${\bm{N}}$ represents the noise distribution $q_{0}$. This formulation includes uniform noise (${\bm{N}}=\frac{1}{d_{Z}}{\bm{1}}{\bm{1}}^{T}$), masking noise (${\bm{N}}={\bm{1}}{\bm{e}}_{\text{Mask}}^{T}$, where ${\bm{e}}_{\text{Mask}}$ is the one-hot vector for the mask label), and marginal noise (${\bm{N}}={\bm{1}}{\bm{m}}^{T}$, where ${\bm{m}}$ encodes the marginal distribution).
Under this formulation, $\bar{{\bm{Q}}}_{t}$ is computed efficiently as:
$\bar{{\bm{Q}}}_{t}=\alpha_{t}{\bm{I}}+(1-\alpha_{t}){\bm{N}},$ (3)
@@
By learning a backward process $p^{\theta}_{s\mid t}(z\mid Z_{t})$ for $s=t+\Delta t$ and sampling iteratively from it, we generate data samples from the noise distribution.
2.3 Temporal Dependencies and Compounding Error
At the beginning of the backward process, $Z_{t}$ contains limited information, resulting in a high-entropy distribution $p^{\theta}_{s\mid t}(z\mid Z_{t})$, from which the elements $z_{s}$ of $Z_{s}$ are independently sampled. At time $s$, the joint distribution from the backward process:
@@ 2.3 Temporal Dependencies and Compounding Error
may deviate significantly from the forward distribution due to early-stage unmasking errors.
In the case of mask diffusion, unmasked elements cannot be masked again, as the conditional distribution collapses to a Dirac delta:
@@
This leads to error accumulation, a phenomenon known as the compounding denoising error (Lezama et al., 2023).
The compounding denoising error is particularly pronounced in mask diffusion. It is somewhat less severe with uniform or marginal noise distributions, as all elements can change at each denoising step, allowing for corrections. However, due to the direct dependence of $z_{s}$ on $z_{t}$, see graphical model of Discrete Diffusion in Figure LABEL:fig:denoising LABEL:fig:denoising_a, the probability of correction is low, and the risk of error propagation and accumulation remains important. Specifically, Gat et al. (2024) have shown that a denoising step in discrete diffusion can be expressed as:
$p_{t+\Delta_{t}|t}(z|Z_{t})=\delta_{z_{t}}(z)+\Delta_{t}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}\left[p_{1|t}(z\mid Z_{t})-\delta_{z_{t}}(z)\right],$ (5)
where $\dot{\alpha_{t}}$ is the $\alpha_{t}$ time-derivative. We interpret the second term on the right-hand side as the probability of modifying an element in a denoising step of size $\Delta_{t}$. Notably, the scaling factor $\Delta_{t}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}$ imposes a strong constraint on the probability of modifying an element. Consequently, the compounding denoising error issue also affects other common noise distributions, such as the uniform and marginal distributions.
To address this issue, works have adopted a discrete version of Predictor-Corrector sampling Campbell et al. (2024); Lezama et al. (2023). In the discrete case, predictor-corrector sampling methods are equivalent to simultaneously performing forward and backward steps. Recently, Gat et al. (2024) generalized the corrector sampling strategy, integrating it in the Flow Matching Framework, and Zhao et al. (2024) proposed informed correctors. Our approach circumvents the need for such correctors or additional sampling steps.
@@ 2.4 Motivation
We argue that the strong dependence on previous time steps in existing models hinders the generative performance of discrete diffusion models for graph generation. These models still struggle to consistently generate graphs with desired structural properties, such as valid molecular graphs or planar graphs.
In this work, we propose a simple yet effective framework that overcomes compounding denoising errors by removing temporal dependencies in the noising process. Our Iterative Denoising approach (Sec. 3) assumes conditionally independent noise distributions over time, alleviating the problem of error accumulation.
@@
Additionally, we introduce a Critic-Guided Sampling procedure (Sec. 4), which prioritizes renoising elements with lower probabilities under the data distribution, further improving generative performance, particularly in molecular graph generation.
3 The Iterative Denoising Model
@@
In this section, we introduce Iterative Denoising (ID), a simple yet effective generative framework for modeling discrete-state graphs. This approach simplifies existing denoising models, such as DDM and DFM, while enhancing generative performance. ID serves as the foundation for Critical Iterative Denoising (CID) which further improves the generative performance.
@@
Similar to DDM, ID consists of both noising and denoising operations. The generative sampling procedure involves iteratively applying small denoising steps, enabling effective graph generation.
@@
Notably, if we have access to a pre-trained DDM or DFM denoiser, i.e., a parameterized model predicting $p^{\theta}_{1|t}(z|Z_{t})$, the ID method is training free. In the following, we present the method in details.

### v2

# Simple and Critical Iterative Denoising: A Recasting of Discrete Diffusion in Graph Generation

@@ Abstract
Discrete Diffusion and Flow Matching models have significantly advanced generative modeling for discrete structures, including graphs. However, the dependencies between intermediate noisy states lead to error accumulation and propagation during the reverse denoising process—a phenomenon known as compounding denoising errors. To address this problem, we propose a novel framework called Simple Iterative Denoising, which simplifies discrete diffusion and circumvents the issue by assuming conditional independence between intermediate states. Additionally, we enhance our model by incorporating a Critic. During generation, the Critic selectively retains or corrupts elements in an instance based on their likelihood under the data distribution. Our empirical evaluations demonstrate that the proposed method significantly outperforms existing discrete diffusion baselines in graph generation tasks.
Machine Learning, ICML, Denoising, Discrete, Diffusion, Iterative, Critic, Critical, Graph, Generative, Model, Flow, Matching, Flow Matching, Compounding errors
@@ 1 Introduction
Denoising models such as Discrete Diffusion and Discrete Flow Matching have significantly advanced generative modeling for discrete structures (Austin et al., 2021; Campbell et al., 2022, 2024; Gat et al., 2024), including graphs (Haefeli et al., 2022; Vignac et al., 2023). Despite their success, these models suffer from a key limitation caused by the dependencies between intermediate noisy states in the noising and denoising processes. Errors introduced early in the process accumulate and propagate, degrading generative performance. This issue is particularly pronounced in mask diffusion, where the noising process progressively masks elements of an instance. The problem has been observed in discrete sequence modeling, where techniques such as corrector sampling have been proposed to mitigate it.
@@
Figure 1: Validity rate (without correction) of generated molecules trained on Zinc250k as a function of the Number of Function Evaluations (NFE) for three models: Discrete Diffusion (baseline), Simple Iterative Denoising (ours), and Critical Iterative Denoising (ours).
@@
In graph modeling, marginal distributions over node and edge attributes have proven to be a more suitable noise distribution than the mask distribution, as corroborated by our experiments. However, we show both theoretically and empirically that compounding denoising errors also affects discrete diffusion and flow matching models when the marginal distributions are used as the noise.
@@
To address this challenge, we introduce a novel framework called Iterative Denoising, which simplifies discrete diffusion by assuming that intermediate noisy states depend only on the clean data, making them conditionally independent of one another. By removing direct dependence on partially denoised instances from previous steps, our framework facilitates error correction and substantially improves generative performance.
@@
Furthermore, we show that our model can be interpreted as selectively corrupting certain elements while leaving others untouched. Building on this insight, we introduce a Critic, which modulates the corruption probability of each element based on the element’s likelihood under the data distribution. We provide theoretical motivation for our approach and demonstrate empirically that the method further enhances models’ performance.
@@
Our empirical evaluations highlight the effectiveness of the proposed framework, demonstrating superior performance over existing discrete diffusion baselines in graph generation tasks. Notably, our method tackles the challenge of generating valid molecular structures. In Figure 1, we see that our method needs only a small number of denoising steps to reach almost 100% molecule validity, while the standard Discrete Diffusion tops at 80%.
Our key contributions are as follows:
  • We identify compounding denoising errors as a key limitation in discrete diffusion across various noise distributions, not just masking, as previously reported.
  • To address this issue, we propose a novel, simple, and effective framework: Simple Iterative Denoising (SID).
  • We extend SID with a novel Critic mechanism that adjusts re-masking probabilities based on the predicted clean sample.
  • Empirically, we show that our approach outperforms standard discrete denoising in graph generation tasks, with the Critic providing additional gains, especially in low-NFE settings.
Our code is available at: github.com/yoboget/sid.
2 Background
@@
Graph diffusion models are powerful methods for generating discrete graph structures, such as molecules. However, existing approaches, including Discrete Diffusion Models and Discrete Flow Matching (DFM), face challenges due to the temporal dependencies of the noising process. These dependencies cause error propagation and accumulation, ultimately degrading generative performance. This section introduces core concepts, discusses these limitations, and motivates our approach.
@@ 2.1 Notation
We define a graph as a set of nodes and edges, denoted by ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$. A graph is represented by its adjacency matrix ${\bm{A}}\in[d_{A}+1]^{n\times n}$, where $n=|{\mathcal{V}}|$, and $[d_{A}+1]=\{1,\dots,d_{A}+1\}$, with $d_{A}$ representing the number of edge types and the additional label corresponding to the absence of an edge. Node attributes, if present, are encoded as integers in an annotation vector ${\bm{x}}\in[d_{X}]^{n}$, where $d_{X}$ is the number of node types.
To facilitate readability and simplify the transferability of the method to other discrete modalities, such as sequence modeling, we use $z$ to denote a single element. In graph modeling, $z$ may refer to either a node or an edge attribute when the formulation applies to both $x^{(i)}$ and $a^{(i,j)}$. The corresponding capital symbol $Z$ represents the entire instance, such as a graph ${\mathcal{G}}$. For simplicity and consistency with common practice, we use $p(x)$ to denote the Probability Mass Function (PMF) $P(X=x)$. The distribution of a data point element is represented by the Dirac delta distribution, denoted as $\delta_{z_{1}}(z)$, with all the mass concentrated on $z_{1}$. We sometimes represent the univariate categorical distribution over the variable $p(z)$ as a vector ${\bm{z}}$, where the $i$-th component $z_{i}$ denotes the probability that $z$ belongs to the category indexed by $i$.
We refer to mask distribution as the distribution in which all the probability mass is concentrated on an additional synthetic attribute labeled as Mask, denoted by $\delta_{\text{Mask}}(z)$. This distribution is sometimes referred to as the absorbing-state distribution (Austin et al., 2021). We define mask diffusion as the class of diffusion models that employ this distribution as noise. We use $q$ to refer to the noising distributions and $p$ to the denoising ones.
@@ 2.2 Discrete Diffusion Models
where $\tau=\{1-\Delta_{t},\dots,t+\Delta_{t},t\}$.
The continuous-time case is described by taking $\lim_{\Delta_{t}\to 0}q_{t|1}(z\mid z_{1})$, which transforms the product into a geometric integral.
For categorical (non-ordinal) discrete data, a popular family of noising distributions is given by:
$q_{t-\Delta_{t}\mid t}(z\mid z_{t})=(1-\beta_{t})\,\delta_{z_{t}}(z)+\beta_{t}\,q_{0}(z).$ (2)
This formulation includes forward processes leading to a uniform noise ($q_{0}(z)=\frac{1}{d_{Z}}\;\forall z$), mask noise ($q_{0}(z)=\delta_{\text{Mask}}(z)$), or a marginal noise ($q_{0}(z)=m_{z}$, where $m_{z}$ corresponds to the empirical proportion of elements with attribute $z$ in the dataset.
Under this formulation, the forward process admits a convenient closed-form solution to compute $q_{t|1}(z\mid z_{1})$ efficiently:
$q_{t|1}(z\mid z_{1})=\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z),$ (3)
@@
This noising process has been expressed using transition matrices in prior works, including Austin et al. (2021), Campbell et al. (2022), and Vignac et al. (2023) for graph-based formulations. We note that standard instantiations of DFM for categorical data with independent coupling (Gat et al., 2024) are equivalent to continuous-time discrete diffusion. Therefore, the statements about discrete diffusion in the following also apply to these standard instantiations of DFM.
By learning a backward process $p^{\theta}_{s\mid t}(z\mid Z_{t})$ for $s=t+\Delta t$ and sampling iteratively from it, we generate data samples starting from the noise distribution.
2.3 Denoising Dependencies and Compounding Error
At the beginning of the discrete diffusion backward process, $Z_{t}$ contains limited information, resulting in a high-entropy distribution $p^{\theta}_{s\mid t}(z\mid Z_{t})$, from which the elements $z_{s}$ of $Z_{s}$ are independently sampled. At time $s$, the joint distribution from the backward process:
@@ 2.3 Denoising Dependencies and Compounding Error
may deviate significantly from the forward distribution due to early-stage denoising errors. This leads to error accumulation, a phenomenon known as the compounding denoising error (Lezama et al., 2023).
The compounding denoising error is particularly evident in mask diffusion, where unmasked elements cannot be masked again, as the conditional distribution collapses to a Dirac delta:
@@
While the issue is particularly pronounced in mask diffusion, we argue that it also affects other noising distributions $q_{0}$ such as the uniform and marginal distributions.
In these cases, the compounding denoising error is somewhat less severe as all elements can change at each denoising step, allowing for corrections. However, due to the direct dependence of $z_{s}$ on $z_{t}$ (red arrow in Figure LABEL:fig:denoising LABEL:fig:denoising_a) the probability of correction is low, and the risk of error propagation and accumulation remains significant. Specifically, Gat et al. (2024) show that a denoising step in discrete diffusion can be expressed as:
$p_{t+\Delta_{t}|t}(z\mid Z_{t})=\delta_{z_{t}}(z)+\\ \Delta_{t}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}\left[p_{1|t}(z\mid Z_{t})-\delta_{z_{t}}(z)\right],$ (5)
where $\dot{\alpha_{t}}$ is the $\alpha_{t}$ time-derivative. The second term on the right-hand side represents the probability of modifying an element in a denoising step of size $\Delta_{t}$. The scaling factor $\Delta_{t}\frac{\dot{\alpha_{t}}}{1-\alpha_{t}}$, in particular, imposes a strong constraint on the probability of modifying an element. Consequently, compounding denoising errors are not limited to mask diffusion, but also affects other common noise distributions, such as uniform and marginal distributions.
@@ 2.4 Motivation
We argue that direct dependence on previous intermediate states in existing models hinders the generative performance of discrete diffusion models for graph generation. These models still struggle to consistently generate graphs with desired structural properties, such as valid molecular graphs or planar graphs.
In this work, we propose a simple yet effective framework that mitigates compounding denoising errors by removing dependencies on previous intermediate states in the noising process. Unlike standard DDM and DFM, which rely on the Markovian assumption, our Iterative Denoising approach (Sec. 3) assumes conditional independence from prior noisy states, alleviating the problem of error accumulation. While our method operates in a discrete setting, it relates to the general degrade-and-restore strategy developed for image generation(Bansal et al., 2023).
@@
Additionally, we introduce a Critic-Guided Sampling procedure (Sec. 4), which prioritizes re-noising elements with lower probabilities under the data distribution, further improving generative performance, especially in molecular graph generation.
2.5 Related works
Our work primarily relates to two areas of research: methods that address compounding denoising errors and generative models for graphs. In the following, we review relevant contributions in both fields.
2.5.1 Addressing Compounding Denoising Errors
Discrete diffusion approaches to mitigating compounding denoising errors follow two main strategies: corrector sampling and planning. Corrector sampling (also known as predictor-corrector sampling) (Campbell et al., 2024; Lezama et al., 2023) involves alternating denoising (backward) steps with noising (forward) steps to improve convergence. Recent work (Gat et al., 2024) integrates denoising and corrector sampling by simultaneously performing $k$ backward and $k-1$ forward steps. In the context of mask diffusion, this method allows for the correction of previously sampled elements. In Section 3, we show that our Simple Iterative Denoising process is formally equivalent to corrector sampling with maximal corrector steps. However, because elements are randomly re-noised, the efficiency of this approach remains limited. Extending this framework, our Critical Iterative Denoising (CID) introduces modulated re-masking probabilities, effectively acting as a probabilistic planner for re-masking decisions. Critically, unlike prior work, CID operates directly on fully denoised predictions.
Planning, as defined by Liu et al. (2025), involves selecting which elements to denoise (i.e., unmask) at each iteration. This selection can be guided by confidence scores from model outputs Zhao et al., 2024b, Zheng et al., 2024, Nie et al., 2025* or by a dedicated planner Liu et al., 2025; Kim et al., 2025* [^1] References marked with ’*’ denote concurrent work, meaning they were published after this article was submitted.. Outside discrete diffusion, confidence-based unmasking has been explored by Chang et al. (2022). Notably, planning does not influence the model’s capacity to revise previously sampled elements. Using mask diffusion, Peng et al. 2025* mitigate this issue by jointly deciding which masked elements to unmask and which unmasked elements to remask. In contrast, our Simple Iterative Denoising approach removes the need to explicitly plan a denoising order by predicting all elements simultaneously at each step. Furthermore, our approach generalizes beyond masking to support various noising distributions.
2.5.2 Graph Generative Models
A key advantage of discrete denoising methods, including discrete diffusion, discrete flow matching, and our iterative denoising, is that they do not depend on a specific element ordering. This property offers a flexible framework for sequence infilling and generation of unordered structures such as sets and graphs. In this work, we focus on the application of denoising models for graph generation.
Recent denoising generative models for graphs leverage the order-agnostic nature of denoising processes to maintain permutation equivariance. Some approaches assume continuous diffusion spaces—such as score-based models (Yang et al., 2019; Jo et al., 2022), diffusion bridges (Jo et al., 2024), and flow matching (Eijkelboom et al., 2024)—but this assumption breaks the discrete nature of graph structures during noising. Discrete diffusion models, whether in discrete time (Haefeli et al., 2022; Vignac et al., 2023) or continuous time (Xu et al., 2024), and discrete flow matching models, have shown strong performance on graph generation. However, these models are affected by compounding denoising errors. Our approach mitigates this issue, leading to improved generative performance. A broader review of graph generative models, including non-equivariant approaches, is provided in Appendix C.
3 Simple Iterative Denoising
@@
In this section, we introduce Simple Iterative Denoising (SID), a simple yet effective generative framework for modeling discrete-state graphs. This approach simplifies existing denoising models, such as DDM and DFM, while enhancing generative performance. SID serves as the foundation for Critical Iterative Denoising (CID), which further improves the generative performance.
@@
Similar to DDM, SID consists of both noising and denoising operations. The generative sampling procedure involves iteratively applying small denoising steps, enabling effective graph generation.
@@
Notably, if we have access to a pre-trained DDM or DFM denoiser, i.e., a parameterized model predicting $p^{\theta}_{1|t}(z\mid Z_{t})$, the SID method is training-free. Table 3 summarizes the key differences between SID, DDM, and DFM. In the following, we present the method in detail.